You are the operations manager of a firm that uses the continuous review inventory control system. Suppose the firm operates 252 days a year and has the following characteristics for its primary item:

   Demand = 25,000 units/year

Ordering cost = $33/order

Holding cost = $4/unit/year

Lead time = 4 days

Standard deviation in daily demand = 3 units

What is the total holding cost per year, including annual holding cost for safety stock (to the nearest whole number)? (Service level=95%)
Answer #1

Annual demand (D) = 25000 units

Number of days per year = 252

Ordering cost(S) = $33

Holding cost(H) = $4

Lead time (L) = 4 days

Standard deviation of daily demand (d) = 3 units

At 95% of service level value of Z = 1.64

Economic order quantity (Q) = √(2DS/H)

= √[(2 X 25000 X 33) /4]

= √(1650000/4)

= √412500

= 642 units

Annual holding cost = (Q/2) H = (642/2) 4 = $1284

Safety stock = Z x d X √L = 1.64 x 3 x √4 = 1.64 x 3 x 2 = 9.84 units

Annual holding cost for safety stock = safety stock x H = 9.84 x 4 = $39.36

Total holding cost per year including annual holding cost for safety stock = Annual holding cost + Annual holding cost for safety stock

= $1284 + $39.36

= $1323.36 or rounded up to $1323